OpenStudy (anonymous):

Sam bought h shirts for $22. He sold them at $6 each. Find how much he earned in terms of h. HELP ME PLEASE :)

OpenStudy (anonymous):

$ he got for selling 'h' shirts= 6*h gain=6*h-22